OpenStudy (anonymous):

How to take integral: (xe power X)dx/Sq of (1+e power X)

zepdrix (zepdrix):

\[\Large \int\limits \frac{x e^x\;dx}{\sqrt{1+e^x}}\quad=\quad \int\limits \frac{e^x\;dx}{\sqrt{1+e^x}}\cdot x\]Hmm from this form I think we can successfully apply `Integration by Parts`. You are familiar with this technique, yes?

zepdrix (zepdrix):

The x isn't outside of the integral. I shouldn't have written it AFTER the dx, sorry that was sloppy of me :)

zepdrix (zepdrix):

\[\Large =\quad \int\limits\limits x\cdot\frac{e^x\;dx}{\sqrt{1+e^x}}\]

zepdrix (zepdrix):

\[\Large\bf u=x,\qquad\qquad dv=\frac{e^x\;dx}{\sqrt{1+e^x}}\] \[\Large\bf du=?, \qquad\qquad v=?\]